OpenStudy (anonymous):

x3 + 5x2 - 11x + 55 = 0

OpenStudy (anonymous):

need help Find the roots of the polynomial equation. -x3 + 5x2 - 11x + 55 = 0 A. , , –5 B. , , 5 C. , –5 D. , , 5

OpenStudy (anonymous):

sorry didnt come out right

OpenStudy (anonymous):

Divide -3x3 - 2x2 - x - 2 by x – 2. A. -3x2 + 4x + 15, R 32 B. -3x2 + 4x + 15 C. -3x2 - 8x - 17 D. -3x2 - 8x - 17, R –36

OpenStudy (raden):

subtitute x = 5 into your equation above, if gives the result equal zero then one of the solution is x = 5

OpenStudy (anonymous):

-X^3+5x^2-11x+55=0 -x^3+5x^2-11x=-55 -x(-x^2-5x+11)=-55 this isn't factorable? Stick it in your calculator and find the intersection, it should be your roots?

OpenStudy (anonymous):

thank you